8. Solomon, Inc., is a rapidly growing chain of commercial banks in north central states. A security analyst's report issued by a national brokerage firm indicates that debt yielding 15%, comprises 25% of Northwest's overall capital structure. Furthermore, both earnings and dividends are expected to grow at a rate of 25% per year.

Currently, common stock in the company is priced at $25, and is not expected to pay dividends during the coming year. This yield compares favorably with the 10% return currently available on risk-free securities and the 16% average for all common stocks, given the company's estimated beta of 2.5.

Calculate Solomon’s component cost of equity using both the capital asset pricing model and the dividend yield plus expected growth model.
Answer #1

Cost of equity using the capital asset pricing model

Cost of equity = Risk-free rate + (Return on all common stocks - Risk-free rate) x Beta factor

Cost of equity = 10 + (16−10) x 2.5

=10+15 = 25

Cost of equity=25%

Cost of equity using the dividend growth plus expected growth model

Cost of equity= (Next year’s dividend / Current price) + Growth rate

= (0 / 25) + 25% = 25%

Cost of equity=25%